OpenStudy (anonymous):

A die is rolled three times. Find the chance that the face with 4 spots appears at least once.

OpenStudy (raden):

by using the complement probability look will becomes easier so that the chance that the face with 4 spots appears at least once is = 1 - P(not 4, not 4, not 4) = 1 - ( 5/6 * 5/6 * 5/6 ) = 1 - 125/216 = 91/216